OpenStudy (anonymous):

Provide a system of TWO equations in slope-intercept form with no solutions. Using complete sentences, explain why this system has no solutions.

OpenStudy (slaaibak):

y = x + 2 y = x+3 x+2 = x + 3 0 != 1 They are parallel lines and will therefore never have a solution.

OpenStudy (anonymous):

thank you very very much!

OpenStudy (anonymous):

THAT'S A GOOD ANSWER SLAAIBAK